there are 7 girls in a class and 3 boys. two students are to be chosen for a special porject. What is the probability both students will be girls?
Answer by stanbon(75887)

there are 7 girls in a class and 3 boys. two students are to be chosen for a special porject. What is the probability both students will be girls?
---
Ans: 7C2/10C2 = 21/45 = 7/15
---------------------------
Cheers,
Stan H.
-----------------------